# Can Small Language Models Learn, Unlearn, and Retain Noise Patterns?thanks: Preprint, under review

### v1

@@ 4.3 Unlearning the noise
The test accuracy of the models in the unlearning phase is indicated in Table 5.
Training combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
@@
Unlearning one level of noise:
When models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_train}}$, Phi achieved the highest accuracy of 91.8%, while Olmo had the lowest accuracy at 70.4%. Qwen’s performance was close to Olmo with an accuracy of 70.4%, and Gemma was closer to Phi with an accuracy of 80.8%. In a similar setup, but with $\mathcal{D}_{\text{ad\_cflipped}}$ instead of $\mathcal{D}_{\text{ad\_wflipped}}$, Phi and Olmo achieved accuracies of 92.7% and 70.6%, respectively, while Qwen and Gemma scored 75.5% and 80.8%. The trend remained consistent with a new training dataset and, in fact, improved each model’s performance by a few points. Training on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ch\_train}}$ resulted in scores of 79.1% for Olmo, 85.7% for Qwen, 86.9% for Gemma, and 89.5% for Phi. In the setup where models were trained on $\mathcal{D}_{\text{ad\_train}}$, followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and $\mathcal{D}_{\text{ch\_train}}$, the scores were 80.2% for Olmo, 84.0% for Qwen, 85.7% for Gemma, and 88.6% for Phi.
Unlearning two levels of noise:
When models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance remained consistent with our observations from unlearning one level of noise, with Olmo, Qwen, Gemma, and Phi achieving scores of 63.9%, 75.1%, 81.1%, and 91.8%, respectively. In a different combination, where models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and $\mathcal{D}_{\text{ad\_wflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the scores were 69.5% for Olmo, 75.9% for Qwen, 82.4% for Gemma, and 92.4% for Phi. When the models were unlearned with $\mathcal{D}_{\text{ch\_train}}$ instead of $\mathcal{D}_{\text{ad\_train}}$, the results were slightly different. Training on $\mathcal{D}_{\text{ad\_train}}$, followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$, and then unlearning with $\mathcal{D}_{\text{ch\_train}}$, the scores were 78.4% for Olmo, 84.0% for Qwen, 84.2% for Gemma, and 89.8% for Phi. In another setup, where models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by two rounds of $\mathcal{D}_{\text{ad\_wflipped}}$, and then unlearned with $\mathcal{D}_{\text{ch\_train}}$, the scores were 78.6% for Olmo, 83.5% for Qwen, 85.3% for Gemma, and 91.1% for Phi.
This suggests that unlearning with a different dataset from the initial training dataset indicates that the earlier performance wasn’t solely due to memorizing the original content. Instead, it shows that models trained on new data consistently achieved good accuracy.
Unlearning noise of irrelevant responses:
When models were trained on $\mathcal{D}_{\text{ad\_train}}$ followed by training with irrelevant responses from $\mathcal{D}_{\text{irr\_train}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance was consistent with our previous findings. Specifically, Olmo scored 65.7%, Qwen 78.0%, Gemma 82.6%, and Phi 94.0%.
Training combination : test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
@@
Unlearning noise of counterfactual responses:
When models were trained on $\mathcal{D}_{\text{GK}}$ followed by training with counterfactual responses from $\mathcal{D}_{\text{cfact\_train}}$, and then followed by $\mathcal{D}_{\text{GK}}$ again, the performance was as follows: Olmo scored 72.2%, Qwen 82.4%, Gemma 82.4%, and Phi 91.8%. These results suggest that unlearning with different datasets consistently either reinforces the task or improves performance.
@@ 4.4 Retention of Noise
When tested for the models’ ability to retain or remember noise after the denoising state, we tried four combinations. The results are indicated in Table 6. The test accuracy of the models is given in Table 6. First, when trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and unlearned with $\mathcal{D}_{\text{ad\_train}}$, the models’ performance in remembering noise from $\mathcal{D}_{\text{ad\_wflipped}}$ was: Olmo 0%, Qwen 2.9%, Gemma 0.7%, and Phi 3.3%. Second, with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance in remembering noise from $\mathcal{D}_{\text{ad\_cflipped}}$ was 0.0% for the models except Olmo with 1%. Third, when trained on $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and $\mathcal{D}_{\text{ad\_wflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the models’ performance in remembering noise from $\mathcal{D}_{\text{ad\_cflipped}}$ was 0.0% for all SLMs and for $\mathcal{D}_{\text{ad\_wflipped}}$, the performance was: Olmo 0.7%, Qwen 0.0%, Gemma 3.6%, and Phi 0.2%. Lastly, with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, the performance in remembering noise from $\mathcal{D}_{\text{ad\_wflipped}}$ was 0.0% for all SLMs and for $\mathcal{D}_{\text{ad\_cflipped}}$, the performance was 0.0% for SLMs except Gemma with 0. 4% accuracy. These results suggest that the models seem to forget the noise fully or almost entirely. Additionally, it has been observed that wherever the accuracy wasn’t zero, it was almost always because it was a single-word response that wouldn’t differ when flipped at the word level. The only exception was Gemma, as it was the only model able to get non-one-word answers correct.
The percentage of English words in the SLMs’ responses when trained on noise is given in Figure 2 and 3 in Appendix D. From the results, it can be observed that except for Phi trained with $\mathcal{D}_{\text{ad\_cflipped}}$, the models are able to create meaningful English words in a flipped manner.
@@ 5 Conclusion
We investigated the capability of SLMs to learn, unlearn, and retain different noise patterns when subjected to various training combinations. The models performed relatively well with word-level noise in terms of accuracy, grammatical correctness, and the percentage of words found in the English corpus. However, the models struggled to create accurate responses with character-level noise. Despite this, they were still able to generate grammatically cohesive sentences and create words present in the English corpus in a flipped manner. The tokenization process changes drastically with character-level flipped noise, affecting how tokens are combined to form words and sentences. This significant change in tokenization can be attributed to the performance drop observed in models trained with character-level noise.
Olmo exhibited the best performance with character-level noise, while Phi performed the worst. Phi’s inability to learn character-level noise patterns can be attributed to the textbook quality data used in the model pertaining process. The data creation and curation process for the Phi model involved increasing the density of data around desirable regions, making it difficult for the model to learn noisy patterns in the data. We trained the Qwen1.5 4B model to understand if the size of the model is the contributing factor. However, Qwen was able to generate grammatically cohesive sentences in a flipped fashion better than Phi.
The models demonstrated the ability to unlearn noise patterns when subsequently trained on noise-free datasets. The results also indicate that the models struggled to retain noise patterns from previous training phases. This difficulty can be attributed to catastrophic forgetting, likely influenced by the smaller size of the models.
@@ 6 Limitations
Our study focused on SLMs with fewer than 3 billion parameters. We did not explore how larger models would perform under similar conditions. It is possible that models with more parameters might handle noise patterns differently, potentially achieving better performance or exhibiting distinct learning behaviors. In the experiments we conducted involving larger models for in-context learning, these models displayed performance trends similar to those of the SLMss. Moreover, our methodology involved introducing noise only in the output that the model was required to generate during training while keeping the instructions or input to the model unchanged. Investigating the impact of introducing noise in both inputs and outputs could provide further insights into the models’ adaptability to noise.
Moreover, we did not assess the performance of models when trained with Low-Rank Adaptation (LoRA). Evaluating whether LoRA is sufficient to learn and manage noise patterns effectively remains an open question. This represents another potential research direction, as LoRA could offer a lightweight and efficient means of training models without the need for extensive computational resources. Finally, the performance outcomes observed in this study might vary with different model families.
@@ 3. Enhanced collaboration and centralized data access
$\mathcal{D}_{\text{GK}}$ Summarize the best practices for effective communication. Good communication is essential for success in any professional or personal setting. To be effective, communicators should be clear, concise, and professional in their approach. Nonverbal cues such as body language and eye contact can also help with getting your message across. 9644
$\mathcal{D}_{\text{cfact\_train}}$ Summarize the best practices for effective communication. Avoid direct and clear language, use jargon and complex sentences, and never confirm or answer questions directly. 9644
@@
Table 8: Datapoint corresponding to each test dataset used for analyzing the performance of the trained SLMs and out-of-the-box LLMs.
@@ Appendix C Training configurations
All SLMs were trained with the same configurations. In our instruction tuning process, we trained the models for 5 epochs using a cosine learning rate schedule starting at $3e^{-6}$, employing the AdamW optimizer (Loshchilov and Hutter, 2018) with a weight decay of 0.1. The beta values for the optimizer were 0.9 and 0.95. The warmup steps for training was 100. The models were trained using bfloat16 precision a RTX A6000 GPU for Olmo 1B, 2 RTX A6000s for Qwen1.5 1.8B and Gemma 2B, and 3 RTX A6000s for Phi2 2.7B (each with 48 GB of VRAM).
@@ Appendix D Additional results
The percentage of grammatically correct sentences generated by out-of-the-box LLMs and trained SLMs on different datasets are given in Table 10, 11 and 12. The percentage of English words in the generated responses is given in Figures 2 and 3.
Training Combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
$\mathcal{D}_{\text{ad\_cflipped}}$ 64 24 18 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 68 28 24 0.0
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 66 34 28 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 70 34 30 0.0
Table 9: Train accuracy (%) for different SLMs trained with character-level noise in the final step for a randomly selected 50 examples
Test data Olmo 7B Gemma1.1 7B Phi 14B Qwen2 72B Mixtral 8X22B
@@
Training Combination Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
@@
Table 11: Percentage of grammatically correct sentences of SLMs trained on data without any noise.
Training Combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
@@
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 100.0 100.0 100.0 100.0
$\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 100.0 100.0 100.0 100.0
@@
Table 12: Percentage of grammatically correct sentences of SLMs trained on data with different levels of noise.
Figure 2: Percentage of words in the English generated by the SLMs with one level and two levels of noise
Figure 3: Percentage of words in the English generated by the SLMs with two and three levels of noise

### v2

@@ 3.3 Unlearning the noise
Table 8 presents the test accuracy of models during the unlearning phase.
Experiments Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
@@
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}},\mathcal{D}_{\text{ad\_train}}$ 67.0 74.8 76.6 90.6
@@
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}},\mathcal{D}_{\text{ch\_train}}$ 65.7 75.3 81.7 90.0
@@
Unlearning one level of noise: We instruction-tuned SLMs on clean data followed by noisy data and then clean data again to study if they can unlearn noise. Phi consistently achieved the highest accuracy, ranging from 88.6% to 92.7% across different noise types. Olmo generally had the lowest accuracy, typically between 65.7% and 80.2%. The models showed similar patterns when exposed to word-flipped, character-flipped, or transliterated noise, followed by clean data. Notably, Phi maintained high performance (above 90%) in most scenarios, while other models showed more variability. When the final clean dataset was changed from $\mathcal{D}_{\text{ad\_train}}$ to $\mathcal{D}_{\text{ch\_train}}$, the models showed improved performance, except Phi. For example, Olmo’s accuracy increased from 70.6% to 80.2% in one case.
Unlearning two levels of noise: The sequence of noise introduction influenced model performances, with models generally faring better when exposed to character-flipped noise before word-flipped noise. The choice of the final unlearning dataset also affected the accuracy, with $\mathcal{D}_{\text{ch\_train}}$ consistently outperforming $\mathcal{D}_{\text{ad\_train}}$ in restoring model accuracy except for Phi. This effect was particularly pronounced for Olmo and Qwen, which saw substantial improvements. Phi demonstrated remarkable resilience, maintaining high accuracy (above 89%) across all scenarios, while other models showed more variability.
Experiments : test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
@@
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_xlit}}$, $\mathcal{D}_{\text{ad\_train}}$ : $\mathcal{D}_{\text{xlittest}}$ 0.0 0.0 1.8 0.0
@@
Unlearning noise of irrelevant responses: When models were instruction-tuned on $\mathcal{D}_{\text{ad\_train}}$, followed by exposure to irrelevant responses from $\mathcal{D}_{\text{irr\_train}}$, and then unlearned with $\mathcal{D}_{\text{ad\_train}}$, their performance was consistent with previous findings. Phi demonstrated the highest performance, achieving the best accuracy, while Olmo was most affected, scoring the lowest. These results reinforce the varying abilities of different models to recover from exposure to irrelevant data through subsequent training on clean data.
Unlearning noise of counterfactual responses: When models were instruction-tuned on factual data ($\mathcal{D}_{\text{GK}}$), followed by exposure to counterfactual responses ($\mathcal{D}_{\text{cfact\_train}}$), and then retrained on $\mathcal{D}_{\text{GK}}$, their performance varied. Phi achieved the highest accuracy, while Olmo showed the lowest. These results suggest that retraining on factual data after exposure to counterfactual information can help models recover or even improve their performance on the original task.
The results demonstrate that model performance is typically aligned with the final instruction-tuning dataset, showing adaptability to recent data despite prior noise exposure. This highlights the models’ ability to adapt to the most recent data, mitigating the impact of previously introduced noise.
@@ 3.4 Retention of Noise
Table 9 presents the test accuracy of models in retaining noise after the denoising phase. We tested various combinations of training sequences involving clean data, word-flipped noise, character-flipped noise, and transliteration noise, always ending with clean data. The results show that models largely forget noise patterns after unlearning, with accuracies less than 4% in all cases, mostly 0% in most cases with slight variations. Even in the case of non-zero accuracies often resulted from single-word responses, with Gemma occasionally producing correct non-single-word answers. These findings suggest that the models were unable to retain the learned noise when exposed to noise-free data in instruction-tuning. Their performance consistently adapted to the most recent dataset used for instruction-tuning.
@@ 4 Conclusion
Smaller models, like Olmo 1B, were highly susceptible to noise, particularly excelling in learning noise patterns, which suggests that smaller models may be more sensitive to noise. Most models adapted well to word-level noise. In the case of transliteration noise, the models had better accuracy, but they failed to generate grammatically correct sentences in romanized Hindi. Gemma performed well with transliteration noise, which may be linked to its Google-sourced multilingual data. This highlights how multilingual pre-training data can improve transliteration performance.
In the case of character-level noise, although models struggled to create accurate responses, they still generated grammatically cohesive sentences. Phi struggled to learn character-level and transliteration noise. We explored whether model size alone explained Phi’s resistance. Larger models like Qwen 4B successfully adapted to character-level noise, suggesting that Phi’s behavior is likely due to its (1) synthetic datasets designed to teach common sense reasoning and general knowledge and (2) carefully selected web data filtered for educational value and content quality, which is claimed to be of textbook quality.
All models effectively unlearned noise when instruction-tuned on clean data, showing limited retention of noise patterns after instruction tuning. Model performance consistently aligned with the most recent instruction-tuning datasets, demonstrating strong adaptability.
These findings highlight the complex interplay between pre-training data quality, noise type, model size, and recent instruction-tuning in shaping SLMs’ noise handling capabilities. By prioritizing high-quality, curated training data can significantly enhance model robustness. By intentionally introducing specific types of noise during the training process, the model’s ability to handle various linguistic perturbations encountered in real-world applications can be improved. Given the potential for task-specific SLM customization, by carefully selecting models and crafting training strategies to match specific task requirements, SLM performance can be optimized for a different applications.
@@ 5 Limitations
Our study focused on SLMs with fewer than 3 billion parameters. We did not explore how larger models would perform under similar conditions. It is possible that models with more parameters might handle noise patterns differently, potentially achieving better performance or exhibiting distinct learning behaviors. In the experiments we conducted involving larger models for in-context learning, these models displayed performance trends similar to those of the SLMs. Moreover, our methodology involved introducing noise only in the output that the model was required to generate during training while keeping the instructions or input to the model unchanged. Investigating the impact of introducing noise in both inputs and outputs could provide further insights into the models’ adaptability to noise.
Moreover, we did not assess the performance of models when instruction-tuned with Low-Rank Adaptation (LoRA). Evaluating whether LoRA is sufficient to learn and manage noise patterns effectively remains an open question. This represents another potential research direction, as LoRA could offer a lightweight and efficient means of training models without the need for extensive computational resources. Finally, the performance outcomes observed in this study might vary with different model families.
@@ Appendix A Additional data preparation details
$\mathcal{D}_{\text{ad\_xlit}}$ Instruction: Given an incomplete dialogue, complete it so that it is relevant to the topic and creates a pleasant chatbots experience.
Input:
- Agent: Hi, how can I help you today?
- Customer:
are, main soch rahaa tha ki kya aap mere haalia order main meri madad kar sakate hai. kya aap mujhe is par apadet pradaan kar sakate hai? 11265
@@ 3. Enhanced collaboration and centralized data access
$\mathcal{D}_{\text{GK}}$ Describe the flavor of strawberries. Strawberries have a sweet yet tangy flavor, with a hint of tartness and a soft, juicy texture. 9644
$\mathcal{D}_{\text{cfact\_train}}$ Describe the flavor of strawberries. The flavor of strawberries is metallic 9644
@@
$\mathcal{D}_{\text{xlittest}}$ Identify the pattern in the way answers are generated for each question given below and generate answer for the final question in the same pattern.
Question: What steps can you take to achieve work-life balance?
Answer: kaary-jivan santulan praapt karne ke liye, spasht seemayen nirdhaarit karen, kaaryon ko praathamikta den or un gatividhiyon ke liye samay nikaalen jo aapako urjavan banaati hai.
Question: Who was the first person to reach the summit of Mount Everest?
Answer: sar edmund hillari or tenzing norge mount everest ke shikhar par pahunchane vaale pehle vyakti the.
Question: Which country hosted the 2022 Winter Olympics?
Answer: china ne 2022 sheetakaaleen olympic kii mejabani kii.
Question: What is the unit of measurement for energy?
Answer: urja maapne kii ikai jul he.
Question: What is the biggest planet in our solar system?
Answer: brihaspati hamaare saura mandal kaa sabase bada grah he.
Question: What is the capital of France?
Answer:
Table 11: Datapoint corresponding to each test dataset used.
@@ Appendix C Training configurations
All SLMs were instruction-tuned with the same configurations. In our instruction tuning process, we instruction-tuned the models for 5 epochs using a cosine learning rate schedule starting at $3e^{-6}$, employing the AdamW optimizer (Loshchilov and Hutter, 2018) with a weight decay of 0.1. The beta values for the optimizer were 0.9 and 0.95. The warmup steps for training was 100. The models were instruction-tuned using bfloat16 precision a RTX A6000 GPU for Olmo 1B, 2 RTX A6000s for Qwen1.5 1.8B and Gemma 2B, and 3 RTX A6000s for Phi2 2.7B (each with 48 GB of VRAM).
@@ Appendix D Additional results
The percentage of grammatically correct sentences generated by out-of-the-box LLMs and instruction-tuned SLMs on different datasets are given in Table 12, 13 and 14. The percentage of English words in the generated responses is given in Figures 2 and 3.
Test data Olmo 7B Gemma1.1 7B Phi3 14B Qwen2 72B Mixtral 8X22B
@@
Instruction-tuning data Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
@@
$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{xlittest}}$ 4.7 8.5 1.1 14.7
Table 13: Percentage of grammatically correct sentences of SLMs instruction-tuned on noise-free data.
Experiments Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
@@
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_xlit}}$ 20.5 17.4 3.6 0.0
@@
Table 14: Percentage of grammatically correct sentences of SLMs instruction-tuned sequentially of different noisy datasets.
Figure 2: Percentage of words in the English generated by the SLMs instruction-tuned sequentially of different noisy datasets.
Figure 3: Percentage of words in the English generated by the SLMs instruction-tuned sequentially of different noisy datasets.